LHC 2012 - latest news

Week 16: Machine coordinators: Jan Uythoven, Ralph Assmann

Plan for the week: Finish intensity ramp-up, stable beams, MD period

 

Date

Summary

Link to report from morning meeting

Monday 23 April

  • 01:00 Start ADT MD
  • 06:00 Start of TS

8:30 meeting

MD1 report

Sunday 22 April

  • 00:14 B2 XPOC and IPOC failed due to no data for the MKD. This was due to the server cfi-ua67-mkdpm2 being down. Reset.
  • 02:30 Aperture MD starting with beam at 4 TeV
  • 07:20 Beam lost at 4 TeV while kicking with AC dipole for aperture measurements
  • 07:20 Preparing RF batch by batch blow-up
  • 09:00 Start RF MD batch by batch blow-up
  • 17:40 Start of BI MD

9:00 meeting

Saturday 21 April

  • 00:33 Interlock during ramp on BLM HV
  • 02:46 Again, interlock BLM HV during ramp, with reduced bunch intensity
  • 05:12 Pilot ramp for tune measurement.
  • 05:26 Dump beams - ramp down - prepare for long. impedance MD
  • 13:30 Start transverse damper (ADT) MD
  • 17:06 Dump at full energy end of ADT MD
  • 18:30 Collimation MD @ 4 TeV starting
  • 20:30 Beam dumped at 4 TeV, collimator moving into beam

9:00 meeting

Friday 20 April

  • 00:40 Injecting physics beam TI8 power converter problem. Piquet.
  • 04:52 Stable beams fill #2536, initial lumi 5.6e33 cm-2s-1
  • 14:40 Programmed dump, end of fill.
  • 15:35 Injecting, but problem on TDI position. Investigate with beam measurements.
  • 23:00 Injection for physics 1380b

8:30 meeting

Thursday 19 April

  • 01:53 RCBV17.R7B2 tripped as happened few days ago: SIS intercepted.
  • 05:54 Stable beams, 1380b, still RF problems B1, abort gap cleaning.
  • 10:09 Dump beam for RF investigations. Access needed for RF.
  • 13:02 Lost communication with QPS on MB.A9L5. Access.
  • 14:22 Access for QPS finished but nor beam from PS until 16:00.
  • 16:03 Injecting probes.
  • 18:23 Stable beams fill #2535, initial 5.4e33 (average). RF OK!
  • 21:00 Start length scale scans for Atlas & CMS, beam unstable: dumped.
  • 21:34 Cryo FIP communication problem S34: mask during injection
  • 22:00 RCBV17.R7B2 tripped again, third time: leave off
  • 23:20 Reset undulator L4, trip, reset, lost cryo condition ML4

8:30 meeting

Wednesday 18 April

  • 02:00 2 x 1380 bunches injected start ramp
  • 02:45 During collision process, vacuum spike on VGPB.219.1R8 at 1.8E-7, valves closure
  • 04:01 Injection kicker timing problems B2, Injection pre-pulse shifted, causing BETS to inhibit beam injection. 144 bunches on TDI.
  • 10:30 MKI resonant charging length increased. Filling 1380 x 1380.          11:48 Fill 1380 x 1380, dumped by operator because losing almost all BPM crates on beam 1 during ramp
  • 13:00 Injecting 1380 x 1380
  • 13:47 Operator dump, Again lost orbit reading on many crates, beam 2
  • 14:30 PS intervention on septum. BPM investigation. Access Atlas and beam dump. Restart of SIS with slightly tighter tolerances on COD trips.
  • 16:32 Injection. BPM investigation ongoing.
  • 17:54 BPM issue solved by rebooting CS-CCR-COLSA (report Alastair Bland).
  • 18:31 1380b filled.
  • 19:19 First stable beams with 1380b. New LHC record luminosity: ~5.1e33.
  • 21:53 Abort gap population creeping up for B1.
  • 22:14 Abort gap cleaning switched on.

8:30 meeting

Tuesday 17 April

  • 01:23 Beams dumped by EIC after some low pile up data for experiments.
  • 03:17 Around 45min of inject and dump in bucket 31181 per both beams.
  • 04:00 Prepare second fill for vdM scans.
  • 06:22 Stable beams, v/d Meer scans for Alice and LHCb.
  • 13:20 End v/d Meer scans.Dump beam.asdf
  • 14:30 Injecting pilot, prepare for abort gap cleaning tests.
  • 15:15 No beam from SPS
  • 16:00 Injecting for abort gap cleaning tests at 4 TeV
  • 19:36 AGC tests finished. Dump beams.
  • 19:40 Ramp down and access for CMS
  • 21:40 End of CMS access - CPS POPS down
  • 22:00 Start injection steering

8:30 meeting

Monday 16 April

        Power converter problem, access needed

        06:30 prepare injection for VdM scans.

        08:17 Injection. Programmed dump at injection, due to low bunch intensities for VdM scan (several bunches at 5e10, which is too close to the BPMSW threshold of 4e10). Alice solenoid tripped (low gas/water pressure in the solenoid), ramping back up. Refill.

        12:44 Ramp fill for vdM scan.

        13:41 Stable beams.

        14:05 vdM scan starting.

        14:21 RCBV17.R7B1 tripped and was caught by the SIS. The orbit was growing and we would have dumped on the excursion otherwise. No error in the PIC history buffer for this circuit.

        At collision settings without beam, Daniel V updated abort gap cleaning parameters.

        14:59 Rampdown and preparing for access for ALICE.

        16:43 Injection.

        17:57 Ramp fill for vdM scan.

        18:34 Stable beams. vdM scans starting.

8:30 meeting